OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = x + 4 and passes through the point (−6, 3).

OpenStudy (anonymous):

Perpendicular lines have a slope that is the negative reciprocal. Then plug it into y = mx + b y = -1x + b 3 = -1(-6) + b 3 = 6 + b b = -3 Therefore, y = x - 3